# On the handling of method failure in comparison studies

### v1

@@ 2.2 Typical manifestations of method failure
Non-convergence and other calculation issues  Authors often encounter the inability of a method to perform a calculation necessary for prediction or estimation, either implied through an error, a warning message, or a method output that is not meaningful. For methods employing an iterative procedure (e.g., an iterative maximum-likelihood procedure), this is often referred to as “non-convergence”, i.e., the method fails to produce a stable and valid output in the pre-specified number of iterations.
@@
Memory issues  Even before exceeding possible time budgets, some methods crash due to memory issues. That is, the given method consumes more memory than is available. Gijsbers et al., (2024) additionally list segmentation faults as memory-related errors.
Runtime issues  A variant of method failure manifested and reported in existing comparison studies is excessive runtime. Note that in some comparison studies, fixed time budgets are predefined (possibly with an additional leniency period) such that exceeding them is declared method failure (Gijsbers et al., , 2024). Other studies do not predefine fixed budgets but find that, due to excessive runtimes, running the entire experiment for the affected method is impossible and must therefore be aborted (Herrmann et al., , 2021). Note that in contrast to the other two manifestations, method failure is actively declared by the authors when runtime issues occur. That is, running the methods under investigation on hundreds or thousands of data sets in the study creates the necessity of time restrictions, which are set by the authors arbitrarily, and without which the method would likely produce a valid output. In particular, in many cases, a corresponding real-world user, running the method on a single data set only, would likely obtain a valid output. This has important implications for the handling of method failure, which we will discuss in Section 3.2.
At first glance, it may seem like calculation, memory, and runtime issues are separate reasons for method failure, such that authors report specific “reasons” for failure for different methods within their study or even different “reasons” for a single method in different instances. For instance, some authors report non-convergence as a “reason” for failure in one method–data set combination and memory issues in another. However, upon closer examination, it becomes evident that these are not so much reasons but rather manifestations of method failure that are just the tip of the iceberg. In essence, method failure (or what might appear as such) is often the result of an individual and complex interplay of many different factors.
For example, Kriegel et al., (2017) show that runtimes can depend on numerous factors: “You might be surprised by the amount of differences in runtime that we point out, spanning up to several orders of magnitude in different implementations of one and the same algorithm. We found major performance differences even within the same toolkit, in different versions or releases: due to bugs, algorithmic improvement (having an indirect impact on the tested algorithm, e.g., in data handling), regressions, and environment changes. Finally, it is also noteworthy that good or bad choice of parameters for different algorithms can lead to void runtime comparisons”.
It is important to emphasize that differences in hardware (e.g., available memory) and/or software (e.g., operating system) can also affect assessment. As an extreme example of software dependency, consider that Herrmann et al., (2021) reported that one of the methods under investigation (sparse group lasso) was completely unusable due to an error in R that occurred using Windows as an operating system, while it did not occur under Ubuntu. As an extreme example of hardware dependency, consider that the feasibility of deep learning to train large artificial neural networks tremendously increases with the use of graphics processing units (GPUs). In essence, this means it is important to keep in mind that a computation that fails or is infeasible in terms of runtime and budget on one specific system setup, may work on a different setup. Note that the reverse situation that a computation that can be successfully conducted on one specific system cannot be conducted—i.e., reproduced—using a (slightly) different setup (e.g., during review), is very common and addressed by the term computational reproducibility.
In addition to these (combinations of) factors associated with a method (or its implementation) and the technical issues mentioned above, failure is typically also related to certain characteristics of the given data. That is, some characteristics can complicate required calculations for a method such that they are completely impossible or cannot be completed within the maximum number of iterations (non-convergence), or they exceed available time or memory budgets. For instance, van Smeden et al., (2016) report non-convergence of logistic regression algorithms when the outcome variable is perfectly separated by one or more predictor variables.
To sum up, oftentimes, certain characteristics of the data cause a certain method, in its current implementation, to fail in a certain computational environment.
2.3 Common approaches to handling method failure
The approaches to handling method failure from the published comparison studies in Section 2.1 can generally be divided into two categories. First, some authors resolve method failure so that no undefined values remain. This includes entirely removing failing methods from the study or modifying its implementation (e.g., adjusting some parameters) to make it work for all data sets. Second, other authors summarize the affected method(s) performance(s) despite undefined performance values. Most commonly, the corresponding data sets are discarded either for all or only for the failing methods. Other authors impute undefined performance values. Again, this is often done in various ways, for example, imputation through the worst possible value (when such a value exists) or through the worst (or mean) value achieved by other methods on the particular data set. Additionally, some authors also report the proportion of (real or simulated) data sets for which each method fails.
In the following Section 3, we discuss why imputation and discarding data (either for all or the failing methods only) are inadequate in most cases of method failure. In particular, we recommend alternative and more appropriate approaches that are based on realistic considerations regarding the research stage in which the study takes place and the corresponding targeted users.
3 Principles and recommendations for handling method failure
3.1 Absolute, relative, unconditional and conditional performance
@@
There is no one-size-fits-all solution to handling method failure, exemplified by the examples in Section 2. This section discusses some principles to guide decisions. To illustrate ideas, we begin with two simple examples and some useful terminology. Suppose a simulation study aims to compare the performance of method A with method B in terms of bias. Method A fails in $20\%$ of repetitions while method B does not fail in any. Alternatively, suppose a benchmark study aims to compare the average accuracy of two methods. Again, method A fails in $20\%$ of benchmark data sets while method B does not fail in any. The hypothetical results for the two cases are given in Table 3(b) (a) and (b).
@@
We contrast the terms unconditional vs. conditional performance and absolute vs. relative performance. Unconditional performance simply means “in all data sets”, while conditional means “in data sets with no method failure”. The unconditional bias of method B is 0.14, while the unconditional bias of method A is not defined. The conditional bias of both methods is 0.09. Parallel descriptions apply to the hypothetical benchmark study in Table 3(b) (b). Absolute performance means “considered on its own” and relative performance means “when compared with another method”. In both Tables 3(b) (a) and 3(b) (b) the relative unconditional performance of methods cannot be quantified because the unconditional performance of method A is undefined. The relative conditional performance can be quantified: in Table 3(b) (a), the two methods have the same bias conditional on non-failure of method A; in Table 3(b) (b), the two methods have the same accuracy conditional on non-failure of method A.
@@
These terms help to distill the problems with method failure in comparison studies. Firstly, the absolute conditional performance of method B should be of little interest. In deciding to use method B, why should we care about its performance when method A, a method we are not going to use, does not fail? We should care about a method’s unconditional performance. The problem is that we cannot compare the methods in terms of their unconditional performance. Comparisons should be performed vertically, or we are not comparing like with like: if we compared the unconditional performance of method B with the conditional performance of method A (which corresponds to discarding data sets for the failing methods), this would reflect poorly on method B, even though both perform equally well for the data sets without failure. This is all the more problematic since method B, in this case, is actually preferable to method A, as it outputs a result for all data sets!
The methods’ conditional performance (“discard corresponding data sets for all methods”) does compare like-with-like. In conducting this comparison, method B’s conditional performance is in both cases (a) and (b) better than its unconditional performance.
3.2 Caution with discarding data and imputation
@@
It is clear that what we would really like to quantify, but cannot, is relative unconditional performance. The superficial solution often proposed when looking at results such as those in Table 3(b) is to consider undefined values due to method failure as missing values and use missing data methods, such as imputation to infer the unconditional performance of method A. After all, results given by different methods tend to be positively correlated, so results given by method B could inform the imputation.
@@
We view it as mistaken to regard method failure as a form of missing data and have been careful to refer to “undefined” rather than “missing” values. The missing data literature defines missing values as those that exist but whose values were not recorded (Carpenter et al., , 2023; Rubin, , 2004). Method failure is far closer to what is called a truncating event in the estimands literature (Kurland et al., , 2009; Kahan et al., , 2024), which denotes a value that does not exist. That is, when a method fails, the result we wished to record simply does not exist for that data set. For this reason, we should be cautious about applying missing data approaches to method failure. In particular, imputing undefined performance values or discarding data sets associated with method failure (either for all methods or the failing methods only) do not enable the comparison of relative unconditional performance. There are two exceptions to this.
First is when a “failure” is recorded but a result could have been obtained; for example, if failure occurs due to excessive runtime but the definition of excessive in the comparison study can safely be judged as shorter than in practice. For cases like this, it may make sense to run the experiment on a randomly selected subset of all simulation repetitions/benchmarking data sets for the method affected by runtime issues (“failure by design”). As this can be seen as missingness mechanism “missing completely at random” (missingness is “unrelated to the observed and unobserved data” (Carpenter et al., , 2023)), discarding the data sets for which the method “fails” or imputing based on the performance results from other methods or data sets may be appropriate. To ensure missingness completely at random, it is important to select the subset of data independently of the data sets for which the method has particularly long runtimes. Researchers should therefore think about possible runtime problems and carry out the corresponding random selection of data sets before conducting the study (see Section 3.3.7).
A second exception is when failure occurs for no reason, such that it is safe to compare the unconditional performance of method B with the conditional performance of method A. However, because method failure always happens for a reason – the interplay of a particular method with a particular data set – we do not believe this would ever occur.
An often more suitable approach to handling method failure than discarding data sets or imputing, which we term the “fallback strategy”, involves specifying a backup method when a failure occurs. This strategy is described in more detail in Section 3.3.5.
Table 4: Recommendations on handling method failure
Caution with Explanation
Imputation In most cases: method failure does not lead to “missing data”
$\rightarrow$ Possible exception: Runtime issues
Discarding data
(for all methods or the
failing methods only)
Usually does not provide assessment and comparison of relative unconditional performance
$\rightarrow$ Possible exception: Runtime issues
Alternative recommendations Explanation
Pinpoint interplay causing failure as the basis of any of the following considerations…
Check for method misuse Is the failing method truly developed for given estimation or prediction task?
Check if data matches scope Are targeted users likely to encounter this kind of data and failure?
Be realistic when modifying method implementations Consider changes in implementations to solve method failure that are realistic for targeted users and research stage. Possibly combine with fallback strategies (see below).
Incorporate fallback strategy Incorporate method pipelines “method A if it succeeds in producing an output, method B otherwise”.
$\rightarrow$ Enables assessment of relative unconditional performance.
Report transparently … - failure proportions, ideally as performance measure. - interplay causing failure with sufficient detail: “Method Z, with an overall failure proportion of $X\%$, fails when … and when … .” - chosen handling(s) and justify choices based on above considerations.
Slimmed-down recommendations
Analyze and report simultaneously:
(i) results after discarding data sets for all methods.
(ii) results after discarding data sets for the failing methods.
(iii) proportions of method failure for all methods.
$\rightarrow$ Report that this does not provide assessment of relative unconditional performance.
3.3 Practical recommendations in the presence of method failure
In the following, we present practical recommendations in the presence of method failure that may be more appropriate than imputing or discarding data sets. We advise implementing these recommendations as comprehensively as possible as they contribute to increased usability of comparison studies to users of the methods. However, we are aware that methodological researchers facing time constraints and page limits may be prevented from following them completely. In Section 3.4, we therefore discuss how a “slimmed-down” version of our recommendations can be applied nevertheless.
3.3.1 Pinpoint interplay of method and data characteristic causing failure
We recommend authors of comparison studies investigate and understand for which types of data sets and why a corresponding method fails, beyond the initially perceived symptom(s). Ideally, they should identify if particular (and if different) data characteristics cause the failures, leading to evaluations such as “method A fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes”. Having identified the dynamic between the given method and data that causes method failure, authors should base their further handling on the considerations and recommendations provided in the following Sections 3.3.23.3.7.
3.3.2 Check for method misuse
When encountering method failure, authors should first check whether the method was truly developed for the given estimation or prediction task, and the type of data (based on the investigations from the previous Section 3.3.1), and if it is used as intended by the developers. If not the case, what appears initially as method failure may actually be method misuse. Method misuse should not be attributed to the method but to the user, that is, the method should not be penalized for failing on a data set or a task it was not intended for. To avoid method misuse and any associated failure in the first place, authors of comparison studies should familiarize themselves thoroughly with the methods under investigation and read the corresponding user manuals carefully. It may also make sense to consult the developers of the individual methods (and possibly involve them in the study).
Note that there may also be cases where a method is commonly misused by real-world users, especially when they are inexperienced. Depending on which kind of user the study aims at (e.g., inexperienced versus experienced), the occurrence of this kind of failure may be relevant to them and should therefore be included in the study. In these cases, the misuse should be clearly emphasized.

### v2

@@ 2.1.1 Common manifestations of method failure
Non-convergence and other calculation issues  Authors often encounter the inability of a method to perform a calculation necessary for prediction or estimation, either implied through an error, a warning, or a meaningless output. For methods employing an iterative procedure (e.g., an iterative maximum-likelihood procedure), this is called “non-convergence,” i.e., the method fails to produce a valid output within the pre-specified number of iterations.
Examples  Zapf et al., (2024) encounter non-convergence for two of the three methods under investigation. One returns an error, while the other outputs the current, non-converged estimate in these instances (Zapf et al., , 2024). Masaoud and Stryhn, (2010) encounter instances of “non-convergence or non-sensible estimates” (Masaoud and Stryhn, , 2010) for some method-data set combinations. In the study by Crowther et al., (2014), the methods are based on numerical integration and encounter estimation difficulties (non-convergence) with “challenging data sets(Crowther et al., , 2014). Hornung et al., (2024) report instances of method failure when methods are “not applicable” (Hornung et al., , 2024) to certain differences between training and test sets regarding the patterns of block-wise missingness. Additionally, they report calculation issues for one method explicitly disallowing certain structures in the training data set, resulting in failure. Calculation issues are reported by Ruxton and Neuhäuser, (2013), where some odds ratio estimation methods fail due to division by 00 when the $2\times 2$ contingency table contains one or more zero-entries. Another case of calculation issues can be found in the simulation study conducted by Dunias et al., (2024). For some data sets, the statistical learning method LASSO does not select any predictor variable, leading to a constant predictor and preventing the assessment of performance regarding calibration slope.
@@
Memory issues  Even before exceeding possible time budgets, some methods crash due to memory issues. That is, the given method consumes more memory than is available.
Examples  Gijsbers et al., (2024) report memory constraint violations, segmentation faults, and Java server crashes as instances of method failure related to memory. Hornung et al., (2024) also encounter memory issues, which they attribute to a specific missingness pattern induced into the training data. Fernández-Delgado et al., (2014) encounter memory issues in combination with large data sets.
Runtime issues  Another form of method failure often reported in comparison studies is excessive runtime. Some studies set fixed time budgets, so exceeding them is considered method failure. Others without fixed budgets encounter method failure when experiments for single methods run for an excessive amount of time (e.g. days), forcing the researcher to abort. Unlike non-convergence and memory issues, method failure is actively declared by the researchers when runtime issues occur, as without these (naturally necessary) limitations, a method might eventually generate an output at some point. Resulting implications are explored in Section 3.1.6.
Examples  While Gijsbers et al., (2024) encounter the excess of predefined time budgets for model training, Herrmann et al., (2021) face runtime issues without fixed budgets. They encounter “computations lasting several days for one single model fit for large data sets” (Herrmann et al., , 2021) for one method.
2.1.2 Common approaches to handling method failure
Based on the published comparison studies, the handling of method failure can be commonly divided into two general approaches. Note that a literature review by Pawel et al., (2024), yielding a structured investigation of the frequencies of handling approaches from published simulation studies, results in four main handlings, which are among those we have commonly observed.
Resolving method failure  Some authors resolve method failure completely so that no instances remain. A straightforward way observable in the literature is to completely remove the affected method from the analysis. For instance, Ruxton and Neuhäuser, (2013) exclude all methods that fail with zeros in the $2\times 2$ contingency table. Masaoud and Stryhn, (2010) use method removal as a last resort in their sequence of handlings when previous measures (see below) have been unsuccessful. This illustrates well that oftentimes, different handlings are applied within the same study, sometimes switching between approaches or applying them sequentially until successful. Note that a more moderate form of method removal can be observed in the benchmark study by Herrmann et al., (2021). For the method encountering runtime issues, they restrict the analysis to the subset of smaller benchmark data sets and report the corresponding results separately from the remaining methods.
An alternative to removing methods is to modify their implementation such that an output can be successfully produced. For instance, in the study by Masaoud and Stryhn, (2010), “different optimization techniques were tried” (Masaoud and Stryhn, , 2010) for one method when the default procedure resulted in failure. Crowther et al., (2014) react to method failure by successively increasing the number of quadrature points until convergence is reached.
Lastly, changing software for the method affected by method failure can also be observed in the literature. Switching the operating system can be observed in the study by Herrmann et al., (2021), who for one method encounter “a fatal error in R under Windows, but not using the Linux distribution Ubuntu 14.04” (Herrmann et al., , 2021). Masaoud and Stryhn, (2010), while remaining in the same operating system, generate the required output for one method in SAS when it fails in R.
Summarizing performance despite method failure Instead of resolving method failure, some authors summarize the methods’ performances despite existing failures. A widespread approach is to discard data sets associated with failure. The first version of this approach, consisting of discarding data sets for the failing methods only, is followed by Zapf et al., (2024) for the method affected by errors due to non-convergence. Fernández-Delgado et al., (2014), comparing the methods’ performance based on two performance measures, discard the associated data sets for the methods affected by method failure in the comparison based on average accuracy. The second version of discarding data, namely discarding the corresponding data sets for all methods, is applied by Hornung et al., (2024). Particularly, they argue against discarding the data sets for the failing methods only by saying that it would not yield a fair comparison since predictive performance is generally associated with the induced pattern of block-wise missingness (Hornung et al., , 2024).
Instead of discarding data, it is also common to impute performance values for the instances of method failure. Note that imputation yields a variety of options. Fernández-Delgado et al., (2014), for instance, impute by the mean accuracy of all remaining methods on the given data set for their second performance measure (Friedman ranking) when a method fails as it requires the same number of performance values across all methods. On the other hand, recall Gijsbers et al., (2024) and their comparison study of AutoML methods in which they impute with a constant predictor when a method fails. They choose this “very penalizing imputation strategy(Gijsbers et al., , 2024), which can be viewed as a form of imputation by the worst possible value, with a similar reasoning to Hornung et al., (2024). That is, failure may be correlated with data set characteristics such that discarding data sets for the failing methods or imputing by the average performance of the given method on the remaining data sets may unfairly favor the failing methods (Gijsbers et al., , 2024). Note that Dunias et al., (2024) also impute by a constant predictor when a statistical learning method fails.
An alternative way of imputing is employed by Bischl et al., (2013), who impute performance values based on the proportion of repetitions in which the affected method fails. Below $20\%$, they impute by sampling from a normal distribution estimated from the remaining performance values of that model. Above $20\%$, on the other hand, they impute by the worst possible value with the reasoning that the method’s behavior is “too unreliable for the current data set” (Bischl et al., , 2013). A similar imputation strategy is employed by Herrmann et al., (2021). While using the same imputation rule when a method fails on more than $20\%$ of data sets, they otherwise impute by the mean performance of the method in the remaining “successful” data sets. In a subsequent project, Nießl et al., (2022) consider alternative procedures, including imputation by the worst possible value and by a value that weights the method’s overall proportion of method failure against its performance in the remaining folds. They show that the approach to imputing the undefined performance values has a notable effect on the performance ranking between the methods.
When summarizing performance despite the failure of an existing method, some authors report the proportion of (real or simulated) data sets for which each method fails. For instance, Hornung et al., (2024) list “the frequency of data sets with missing results” (Hornung et al., , 2024) for the methods affected by method failure. Gijsbers et al., (2024) provide multiple failure plots, consisting of the number of failures per method with differentiation of the different failure types, the number of failures by size of the associated data set, and boxplots of the training durations for each method with indicators of excessive runtimes. Zapf et al., (2024) even include convergence as a performance measure, supported by a table of the number of converged simulation runs for each simulation scenario across the methods in the supplement.
2.2 Why popular handlings are often inadequate
2.2.1 Why discarding data is inadequate
@@
Currently, there is no one-size-fits-all solution to handling method failure, exemplified by the examples in Section 2. This section discusses some principles to guide decisions. To illustrate ideas, we begin with two simple examples and some useful terminology. Suppose a simulation study aims to compare the performance of method A with method B in terms of bias. Method A fails in $20\%$ of repetitions while method B does not fail in any. Alternatively, suppose a benchmark study aims to compare the average accuracy of two methods. Again, method A fails in $20\%$ of benchmark data sets while method B does not fail in any. The hypothetical results for the two cases are given in Table 3(b) (a) and (b).
@@
We contrast the terms unconditional vs. conditional performance and absolute vs. relative performance. Unconditional performance simply means “in all data sets,” while conditional means “in data sets where this method did not fail.” The unconditional bias of method B is 0.14, while the unconditional bias of method A is not defined. The conditional bias of both methods is 0.09. Parallel descriptions apply to the hypothetical benchmark study in Table 3(b) (b). Absolute performance means “considered on its own” and relative performance means “when compared with another method.”
These terms help to distill the problems with method failure in comparison studies. Clearly, we want to compare the methods’ unconditional performances. However, in both Tables 3(b) (a) and 3(b) (b), the relative unconditional performance of methods cannot be compared because the unconditional performance of method A is undefined. In contrast, the relative conditional performance can be quantified and compared (“discarding data sets associated with failure for all methods”): in Table 3(b) (a), the two methods have the same bias conditional on non-failure of method A; in Table 3(b) (b), the two methods have the same accuracy conditional on non-failure of method A. However, the conditional performance of method B should be of little interest. Why should a method’s performance depend on the behavior of another method regarding failure? In Table 3(b), method B’s performance improves when assessed based on only those data sets where A does not fail. This becomes even more complex as additional methods are included in the comparison, since discarding the corresponding data sets for all methods requires conditioning on the non-failure of each one.
@@
Alternatively, one could intuitively proceed to compare the unconditional performance of method B with the conditional performance of method A (“discard data sets for the failing methods only”). However, comparisons should be performed vertically, or we are not comparing like with like: In the context of Table 3(b), this would reflect poorly on method B, even though both perform equally well for the data sets without failure. This is all the more problematic since method B, in this case, is actually preferable to method A, as it outputs a result for all data sets! Again, this becomes even more pronounced as additional methods are added to the comparison, since different methods typically fail on different datasets (see Section 3.1.1).
2.2.2 Why imputation is inadequate
@@
Another superficial solution often proposed when looking at results such as those in Table 3(b) is to interpret method failure as causing missing values and use missing data methods, such as imputation, to infer the unconditional performance of method A. After all, results given by different methods tend to be positively correlated, so results given by method B could inform the imputation (analogously, it may be intuitive to impute with the results of method A on the remaining data sets as results tend to be correlated).
@@
We view it as mistaken to regard method failure as a form of missing data and have been careful to refer to “undefined” rather than “missing” values. The missing data literature defines missing values as those that exist and were intended to be recorded but for one reason or another are “not observed” (Rubin, , 2004; Carpenter et al., , 2023). Method failure is far closer to what is called a truncating event in the estimands literature (Kurland et al., , 2009; Kahan et al., , 2024), which denotes a value that does not exist. That is, when a method fails, the result we wished to record simply does not exist for that data set. Method failure should, therefore, not be viewed as obscuring the method’s results on a given data set. Instead, failure is the result we get! For this reason, we should be cautious about applying missing data approaches and particularly imputation to method failure. Note that this also includes measures of “weighted performance,” which weight the performance of a method according to the proportion of data sets with failure for that method. Using the correspondence between weighting and imputation approaches, this implies that weighting implicitly assigns some performance value in instances of method failure. The consequence of our dissuasion from “missing data-inspired” approaches is that our work focuses intensively on alternative ways to deal with method failure.
Note that there is an exception to method failure not leading to “missing” data. This is when a “failure” is recorded but a result could have been obtained; for example, if failure occurs due to excessive runtime but the definition of excessive in the comparison study can safely be judged as shorter than in practice. We refer to this case in Section 3.1.6.
3 Principles and recommendations for handling method failure
The following section presents guidelines for handling method failure that are more appropriate than discarding data or imputation. Since, as previously mentioned, method failure can only be directly observed after the execution stage, the order of our recommendations aligns with the natural process of investigating these failures post-execution. However, we also highlight key considerations that can be addressed during the planning stage to help reduce both the potential for and unpredictability of failure in the first place.
Using the terminology of Heinze et al., (2024), we distinguish between comparison studies in “early” and “late” research phases. Early phase studies investigate if a given method can be “used with caution” in settings similar or only slightly different from its original target setting, possibly refining or extending it while also making limited comparisons. Late phase studies, however, focus on realistic comparisons across diverse scenarios to guide real-world users. We make differentiated recommendations when necessary.
3.1 Practical recommendations for the handling of method failure
3.1.1 Consider method failure as the result of a complex interplay
When looking beyond the manifestations of method failure summarized in Section 2.1.1, it becomes evident that these are just the tip of the iceberg. In the end, regardless of how method failure manifests, it is rooted in an individual and complex interplay of interconnected factors, that is
certain data set characteristics cause the method to fail in its current implementation (and on the given hardware and software).
Besides influencing runtime and memory consumption (Zimmermann, , 2020), data set characteristics contributing to method failure are highlighted by the observation that a method usually fails for specific data sets, not all of them. Certain data characteristics can complicate or completely prevent required calculations of a method within the maximum number of iterations (non-convergence). Recall the study of van Smeden et al., (2016) addressing perfect separation of the outcome variable by one or more predictor variables as a cause for non-convergence of logistic regression algorithms. Additionally, Fernández-Delgado et al., (2014), in their extensive comparison study of supervised classification methods, list “collinearity of data, singular covariance matrices, and equal [predictor values] for all the training [observations], […] discrete predictor variables, classes with low populations, or too few classes” (Fernández-Delgado et al., , 2014) as data set characteristics associated with method failure.
Conversely, the association of method failure with the (implementation of) a given method becomes clear from the usual observation that not all methods fail for the same data sets. That is, methods often vary in their ability to handle specific data set characteristics. For instance, in our illustrative study in Section 4.1, some odds ratio estimation methods are equipped with a way around “zeros,” while others fail.
In addition to the interplay of method and data set characteristics, however, software can also affect the (non-)failure of a method such that results are obtained in one software but not another. Consider, for instance, that Herrmann et al., (2021) reported that one of the methods under investigation was completely unusable due to an error in R that occurred using Windows as an operating system, while it did not occur under Ubuntu. The dependency on hardware, on the other hand, becomes clear in how working memory, besides being directly associated with possible memory issues, can also considerably affect runtimes.
Ultimately, the core of method failure–stemming from the interplay of data set, method, software, and hardware–is considerably more informative for addressing method failure than the manifestation of failure itself. We, therefore, recommend that authors of comparison studies investigate and understand the interplay underlying failure for each instance after the execution phase. Ideally, they should identify if particular (and if different) data characteristics cause the failures, leading to evaluations such as “method A fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes.” As a preparatory measure in the planning stage, authors should check user manuals and help pages for any described problems related to certain data structures for all methods under investigation.
3.1.2 Check for the correct use of the method
When encountering method failure in the execution stage, authors should first check whether the method was truly developed for the given estimation or prediction task and the type of data (based on the investigations from the previous Section 3.1.1), and if it is used as intended by the developers. If this is not the case, a distinction should be made between two scenarios. First, a method in the study is used in a slightly different (but not contraindicated) context than originally intended by its developers. We call this “use beyond the original scope.” Including this method in the study may be justified when real users commonly use this method “beyond its original scope,” however, authors should make a corresponding remark in their manuscript. Second, if using the method in the given context is explicitly contraindicated by the method developers, including it in the study would correspond to method misuse. A method should not be penalized for failing on a data set or a task it was explicitly not intended for. Even if commonly (but probably unknowingly) “misused” by real users, this method should be excluded from the study, and the contraindication should be stated clearly in the manuscript. If none of the scenarios apply, authors should resort to Section 3.1.3.
All assessments regarding the use of a given method should be based on the familiarization with all methods from the experiment, which should already take place in the planning stage. This underlines the importance of carefully reading corresponding user manuals and checking for publications from real-world applications in which these methods are commonly used. This aspect should also be checked when planning the data generation in simulation studies and selecting benchmark data for benchmark studies. When in doubt, authors may also consult the developers of the individual methods (and possibly involve them in the study).